OpenStudy (anonymous):

okay i just need process for this to find root using bisection method. x^2-4x-10=0

OpenStudy (anonymous):

\[\Delta=b^2-4ac=16-4(-10)(1)=16+40=56\\ x=\frac{-b\pm\sqrt{\Delta}}{2a} \]

OpenStudy (anonymous):

no no no bisection method

OpenStudy (anonymous):

ooh k 1. find [a,b] where f(a) and f(b) will have opposite signs

OpenStudy (anonymous):

@Eskijoe no

OpenStudy (anonymous):

2. c=(a+b)/2 3. find f(c) 4. see if [a,c] or [c,b] have opposite signs 5. proceed with step 2 with the new interval

OpenStudy (anonymous):

@electrokid how do i find that? f(a) f(b) with opposite signs

OpenStudy (anonymous):

@Eskijoe why not? the discriminant is +ve it has real roots

OpenStudy (anonymous):

@aajugdar if the interval is not provided, start with a +ve x and -ve x by trial-error

OpenStudy (anonymous):

oh okay got ya

OpenStudy (anonymous):

lets take a=0. f(a)=-10<0 now, we start our hunt for b such that f(b)>0

OpenStudy (anonymous):

hint: x(x-4)=10

OpenStudy (anonymous):

oh sorry misread -10 as 10, lol

OpenStudy (anonymous):

here f(0) = -10 f(-2) = 2

OpenStudy (anonymous):

we need x(x-4)>10

OpenStudy (anonymous):

b=6 should satisfy

OpenStudy (anonymous):

10? you said opposite signs ryt

OpenStudy (anonymous):

iter#1 [a,b]=[0,6] -> [-,+] c = 3 f(c)=9-12-10<0 iter#2 [a,b]=[3,6] -> [-,+] c=(3+6)/2=4.5 f(c).... and continue till you get f(c) close to 0 to your tolerance value

OpenStudy (anonymous):

f(0) = -10 and f(2)= 2?

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

f(-2) i meant

OpenStudy (anonymous):

remember, interval [a,b], a<b

OpenStudy (anonymous):

yes you can have [-2,0] its perfectly fine

OpenStudy (anonymous):

so i can say root lies btwn 0 and -2?

OpenStudy (anonymous):

remember.. this is a quadratic function and has 2 distinct roots. we just found the two intervals where they are

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

-2+0/2 = -1 f(-1) = -5 so its like (-2,-1) (-1,0) now

OpenStudy (anonymous):

how do i proceed from here

OpenStudy (anonymous):

no no f(-2)>0 f(-1)<0 f(0)<0 so, opposite signs.. new interval would be [-2,-1]

OpenStudy (anonymous):

oh got ya

OpenStudy (anonymous):

you have just narrowed down your domain. we keep narrowing down and get closer to the thief based on the sign clues

OpenStudy (anonymous):

that is bisection method, no hint how much closer in regula falsi, we use slope of "secant" as clue.. a linear hint of closeness in Newton raphson, we use slope of "tangent" as clue... quadratic hint of closeness

OpenStudy (anonymous):

you can do this really quick in excel too..

OpenStudy (anonymous):

thanks :) got it